what is the solution to each system of equations ? if there is no solution , explain ...
A. 3x + y = 20
x = 4 = y
B. (y+2)=1/2(x-8)
(y-1)=1/2(x+4)

Answer by Alan3354(69443)

A. 3x + y = 20
x = 4 = y *** Too many equal signs
----------------------
B. (y+2)=1/2(x-8) times 2 --> 2y+4 = x-8
(y-1)=1/2(x+4) times 2 --> 2y-2 = x+4
========================
2y+4 = x-8
2y-2 = x+4
----------------- Subtract
6 = -12
Inconsistent, no solution.